OpenStudy (anonymous):

Prove that if 2^n − 1 is prime, then n is prime for n being a natural number.

OpenStudy (anonymous):

can't be factored

OpenStudy (anonymous):

http://primes.utm.edu/notes/proofs/Theorem2.html "Let r and s be positive integers, then the polynomial xrs-1 is xs-1 times xs(r-1) + xs(r-2) + ... + xs + 1. So if n is composite (say r.s with 1<s<n), then 2n-1 is also composite (because it is divisible by 2s-1)."

OpenStudy (anonymous):

Let n = ab (ie composite) 2^ab-1 = (2^a-1)(1+2^a +2^(2a)+ .....2^(a(b-1)) and a can be swapped for b as well. Since u can factor when n = ab, n must be prime for 2^n-1 to be prime.

OpenStudy (anonymous):

i don't get what linjaaho is trying to say. it's stated that 2^(n) -1 is prime and not composite. and why does the proof involves 2^(n) -1 to be composite?

OpenStudy (anonymous):

estudier. i get this part "2^ab-1 = (2^a-1)(1+2^a +2^(2a)+ .....2^(a(b-1)) and a can be swapped for b as well. " but i'm lost after that.

OpenStudy (anonymous):

It's saying that if n is composite then 2^ab-1 is not prime, so for it to be prime n must be prime.

OpenStudy (anonymous):

thanks